OpenStudy (kalli):

Can someone help me with this problem Solve each equation on the interval [0, 2pi]. 4sin^2-3=0 is the equation

OpenStudy (kalli):

Actually, I think I may have got it but can someone check my answers?

OpenStudy (cwrw238):

what do you get?

OpenStudy (cwrw238):

first what do you get for sin x?

OpenStudy (kalli):

4sin^2x - 3 = 0 sin^2x = 3/4 sinx = √3/2 or sinx = -√3/2

OpenStudy (cwrw238):

yes correct

OpenStudy (kalli):

So would my answer be x= pi/3 or 2pi/3 ?

OpenStudy (cwrw238):

it asks for values in the interval 0 to 2pi and the sine can be negative or positive so there is an angle in each quadrant which satisfies the equation.

OpenStudy (kalli):

So I should have 4 answers then right?

OpenStudy (cwrw238):

pi/3 and 2pi/3 are 2 answers

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):